Q2. A pulsed laser has an average output power of 12W. Each pulse consists of light at wavelength 400nm and has 16ms duration. Compute the number of photons that are emitted in a single pulse?
1
Expert's answer
2020-06-08T10:16:13-0400

The energy of each photon with wavelength "\lambda":


"E_1=h\nu=h\frac{c}{\lambda}."


The electrical energy required to provide 12 W during 16 ms:


"E=Pt."

This energy will be enough for N photons:


"Pt=N\cdot h\frac{c}{\lambda},\\
\space\\
N=\frac{Pt\lambda}{hc}=\frac{12(16\cdot10^{-3})(400\cdot10^{-9})}{(6.626\cdot10^{-34})(3\cdot10^8)}=3.87\cdot10^{17}."